# Driving by Hybrid Navigation: An Online HD-SD Map Association Framework and Benchmark for Autonomous Vehicles

### v2

@@ 5.1 Overall Architecture
Figure 5: Overview of Map Association Transformer.
@@
As depicted in Fig. 5, the Map Association Transformer (MAT) is a transformer specifically designed for map association. All inputs are vectorized representations $\mathcal{V}=\{\vec{v}_{1},\vec{v}_{2},\ldots,\vec{v}_{N}\}$, where each vector $\vec{v}_{i}$ is parameterized by two endpoints and direction: $\vec{v}_{i}=[p_{i1}^{x},p_{i1}^{y},p_{i2}^{x},p_{i2}^{y},\theta_{i}]$, with $\theta_{i}=\arctan\left(\frac{p_{i2}^{x}-p_{i1}^{x}}{p_{i2}^{y}-p_{i1}^{y}}\right)$ and $p_{i1},p_{i2}\in\mathbb{R}^{2}$ being the start/end points. The input maps are composed of an SD map, an HD map, and a boundary. The SD map ($\mathcal{G}_{\mathcal{R}}$) comprises road vectors =$\mathcal{R}=\{r_{1},\ldots,r_{m_{r}}\}$, which form a graph with topological edges $\mathcal{E}_{\mathcal{R}}$. Each road $r_{j}$ is transformed into an ordered sequence of vectors $\mathcal{R}_{j}=\overrightarrow{q_{j1}q_{j2}},\overrightarrow{q_{j2}q_{j3}},\ldots$ through its parameterized segments. The HD map ($\mathcal{G}_{\mathcal{L}}$) consists of centerline vectors =$\mathcal{L}=\{l_{1},\ldots,l_{m_{l}}\}$ that represent the centerlines. Each centerline $l_{i}$ is transformed into vectors $\mathcal{L}_{i}=\overrightarrow{p_{i}^{1}p_{i}^{2}},\overrightarrow{p_{i}^{2}p_{i}^{3}},\ldots$ based on consecutive points $p_{i}^{j}$. The boundary ($\mathcal{B}$) includes the boundary vectors =$\mathcal{B}=\{b_{1},\ldots,b_{m_{b}}\}$, converted similarly to the roads: $b_{j}\to\mathcal{B}_{j}=\overrightarrow{h_{j1},h_{j2}},\overrightarrow{h_{j2}h_{j3}},\ldots$. These vectors are processed by the vector embedding module, which maps each 5D vector $\vec{v}_{i}$ to a high-dimensional feature $F_{\vec{v}_{i}}\in\mathbb{R}^{C}$ via a two-layer MLP. The outputs are aggregated into feature matrices: $F_{road}\in\mathbb{R}^{N_{r}\times C}$, $F_{centerline}\in\mathbb{R}^{N_{l}\times C}$, and $F_{boundary}\in\mathbb{R}^{N_{b}\times C}$, where $N_{(\cdot)}$ denotes the total number of vectors (e.g. $N_{r}=\sum_{j=1}^{m_{r}}\text{length}(\mathcal{R}_{j})$).
@@
Subsequently, the vector tokens are input into a transformer network. MAT consists of stacked MAT blocks to extract hierarchical features, each block containing Path-Aware Attention (PA), Spatial Attention (SA), and feed-forward network (FFN). In the association head, each road token requires the pooling to obtain a representative token $\bar{F}_{road}$ for the current road. The association of the centerline with the road is calculated by combining attention between $\bar{F}_{road}$ and $F_{centerline}$, generating the probability distribution of the associations of the SD-HD map. The association probabilities are further refined by a post-processing method to enforce topological constraints. The details of the implementation are provided in supplemental material.
@@ 5.2 Path-aware Attention
Attention on path order. Inspired by MapTRv2 [17], we introduce the path-aware attention to iteratively refine the global features of the tokens within the road / centerline graphs. Unlike MapTRv2, which focuses on local path segments, we explicitly model long-range dependencies by constructing paths from root nodes to leaf nodes in the graph. This ensures accurate association across distant vectors, albeit at the cost of redundant computations for overlapping path segments.
In path-aware attention, we first reorder vector tokens to align with their path indices: Given a path token sequence $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{k}\}$, the vector tokens are rearranged into $\mathcal{V}_{\text{path}}=\{v_{i_{1}},v_{i_{2}},\ldots,v_{i_{m}}\}$, where each $v_{i}$ belongs to a sub-path in $\mathcal{P}$. After computing attention over $\mathcal{V}_{\text{path}}$, we apply a scatter-mean operation to project tokens back to the original instance order:
$v_{j}^{\text{inst}}=\frac{1}{|\mathcal{I}_{j}|}\sum_{i\in\mathcal{I}_{j}}v_{i}^{\text{path}},\quad\text{where }\mathcal{I}_{j}=\{i\mid\text{token }v_{i}^{\text{path}}\text{ maps to }v_{j}^{\text{inst}}\}.$ (4)
RoPE in PA. We apply multi-dimensional Rotary Position Embedding (RoPE) [28, 2] to encode positional information in path-aware attention. Specifically, RoPE operates on two dimensions: path-level order, which represents the ordinal position of a vector within its parent path (e.g. $\text{pos}_{\text{path}}\in\mathbb{Z}$), and instance-level order, which denotes the ordinal position of the vector within its local segment (e.g. $\text{pos}_{\text{seg}}\in\mathbb{Z}$). This dual-axis encoding scheme enables the model to disentangle global path structures from local geometric details by explicitly modeling both hierarchical relationships and sequential dependencies among vectors.
@@ 5.3 Spatial Attention
Overlook of spatial. Existing transformer models for map processing often neglect explicit spatial coordinates due to their generative nature, making it difficult to associate tokens with physical locations. In contrast, our task benefits from explicit spatial annotations (e.g., GPS coordinates), enabling precise geometric reasoning.
Attention with Vector Serialization. Drawing inspiration from PTv3 [31], we propose a spatial attention mechanism based on vector serialization. Each vector token $\vec{v}_{i}=(p^{1},p^{2})\in\mathbb{R}^{2}$ is encoded in a 3D spatial coordinate $(x,y,r)$, where $x=\lfloor\frac{p^{1}_{x}+p^{2}_{x}}{2g}\rfloor$ and $y=\lfloor\frac{p^{1}_{y}+p^{2}_{y}}{2g}\rfloor$ represent the integer grid coordinates of the vector’s centroid ($g=0.1$m), and $r=\left\lfloor\frac{\theta_{i}}{2\pi/R}\right\rfloor$ denotes the quantized direction angle $\theta_{i}=\arctan\left(\frac{p^{2}_{y}-p^{1}_{y}}{p^{2}_{x}-p^{1}_{x}}\right)$ divided into $R=16$ segments. These coordinates are then assigned to a 1D sequence via a space filling curve $\varphi^{-1}:\mathbb{Z}^{3}\to\mathbb{Z}$, such as the Hilbert curve [11] and the Z curve [24], which preserves spatial location by ordering vectors based on their geometric proximity.
After serialization, tokens are grouped by receptive field size and processed via grouped attention (similar to PTv3). Finally, an inverse serialization operation restores the original token order.
RoPE in SA. In spatial attention, multi-dimensional RoPE encodes absolute spatial positions using $(x,y,r)$, enhancing the model’s ability to focus on geospatial relationships. This complements the relative position bias in standard self-attention.
@@ 5.4 Association and Loss Function
Association. The association between the roads and the centerline is calculated through a cross-attention mechanism. For each road $j$, we first aggregate its token features $\{F^{road}_{j1},\ldots,F^{road}_{jN}\}$ into a representative feature $\bar{F}^{road}_{j}=\frac{1}{N}\sum_{n=1}^{N}F^{road}_{jn}$, where $N$ denotes the number of road tokens on roads $r_{j}$ and $F^{road}_{jn}\in\mathbb{R}^{d}$. The association probability $Prob_{ij}$ between centerline $i$ and road $j$ is then calculated as:
$Prob_{ij}=\frac{\exp\left(\frac{F^{cl}_{i}\cdot\bar{F}^{road}_{j}}{\sqrt{d}}\right)}{\sum_{k=1}^{K}\exp\left(\frac{F^{cl}_{i}\cdot\bar{F}^{road}_{k}}{\sqrt{d}}\right)},$ (5)
where $F^{cl}_{i}\in\mathbb{R}^{d}$ is the centerline token feature, $d$ is the feature dimension, and $K$ represents the total number of road. This formulation normalizes the similarity scores in all roads for each centerline $i$, ensuring a valid probability distribution.
Loss Function. We optimize the model using a combination of Cross-Entropy (CE) Loss and Connectionist Temporal Classification (CTC) Loss. The CE Loss monitors the classification of the centerline by maximizing the log-likelihood of ground-truth associations. To enhance topological consistency, we further apply CTC Loss to align centerline token sequences with road structures. Specifically, each centerline token sequence $\{F^{cl}_{i1},\ldots,F^{cl}_{iT}\}$ is treated as a temporal signal, and the CTC loss enforces alignment with the road token sequence $\{F^{road}_{j1},\ldots,F^{road}_{jT}\}$. The total loss is a weighted sum:
@@ 5.5 Topology Post-Process
We formalize topological decoding as a structured prediction on all path of the centerline $\mathcal{P}_{j},\ j\in[1,\cdots,K]$. $K$ is the number of total paths. The two-stage decoding process operates as follows:
@@
Topological-constraint Beam Search. Based on beem search, topological-constraint beam search makes the following two improvements:
  • Modify the one-way search to implement a bidirectional search starting on $T_{max}$.
  • When generating new predictions, instead of using the approach of taking the maximum value from all roads, we decode under the constraint of connectivity provided in the road network $\mathcal{E}_{r}$, thus ensuring that the connectivity of the road sequence corresponding to the lane path in the decoding result is consistent with the representation of the road network.
Detailed expressions of the topological-constraints Beam Search, including formula descriptions, are included in supplemental material.
@@ 6 Experiment
Methods A-F150 A-F175 A-F195 A-P50:95 A-R50:95 A-F150:95 La./ms
KNN [5] 36.8 36.4 32.4 47.9 27.2 34.6 313
HMM [25] 37.6 36.9 32.6 49.6 28.3 36.0 561
MAT(Ours) 43.4 43.0 38.3 56.1 33.9 42.3 74
Table 1: Result on OMA. La. means latency.
      Methods       A-P50       A-P75       A-P95       A-P50:95       La./ms
      KNN [5]       72.2       69.9       59.4       68.6       299
      HMM [25]       73.8       71.5       60.3       70.1       465
      MAT(Ours)       81.6       80.4       69.0       78.7       70
Table 2: Result on OMA-GT. La. means latency.
6.1 Implement detailed
We train our models from the beginning for a total of 50 epochs utilizing the AdamW optimizer. A cosine-decay learning rate scheduler is employed, incorporating a linear warm-up phase lasting for one epoch. The starting learning rate, weight decay, and batch size are set at 0.0001, 0.05, and 128, respectively, using a NVIDIA A6000 GPU. The latency of MAT operates on an NVIDIA A6000 with pytorch, while KNN and HMM function on an Intel(R) Xeon(R) Platinum 8369B. The KNN and HMM code is developed using Python and Numpy without optimization. In practical applications, highly optimized C++ code could provide better time efficiency.

### v3

@@ 5.1 Overall Architecture
Figure 5: Overview of Map Association Transformer (MAT). The framework processes vectorized roads in SD map and centerlines/boundaries in OP map through NN stacked layers containing Spatial Attention (for global context via curve-based serialization) and Path-Aware Attention (for topological alignment via path indexing). The Association Head then aggregates road features and calculates association probabilities with centerline tokens to generate the final navigation refinement result.
@@
As depicted in Fig. 5, the Map Association Transformer (MAT) is a transformer specifically designed for real-time map association. All inputs are vectorized representations $\mathcal{V}=\{\vec{v}_{1},\vec{v}_{2},\ldots,\vec{v}_{N}\}$, where each vector $\vec{v}_{i}$ is parameterized by two endpoints and direction: $\vec{v}_{i}=[p_{i1}^{x},p_{i1}^{y},p_{i2}^{x},p_{i2}^{y},\theta_{i}]$, with $\theta_{i}=\arctan\left(\frac{p_{i2}^{x}-p_{i1}^{x}}{p_{i2}^{y}-p_{i1}^{y}}\right)$ and $p_{i1},p_{i2}\in\mathbb{R}^{2}$ being the start/end points. The input maps are composed of an SD map, an OP map, and a boundary. The SD map ($\mathcal{G}_{\mathcal{R}}$) comprises road vectors $\mathcal{R}=\{r_{1},\ldots,r_{m_{r}}\}$, which form a graph with topological edges $\mathcal{E}_{\mathcal{R}}$. Each road $r_{j}$ is transformed into an ordered sequence of vectors $\mathcal{R}_{j}=\overrightarrow{q_{j1}q_{j2}},\overrightarrow{q_{j2}q_{j3}},\ldots$ through its parameterized segments. The OP map ($\mathcal{G}_{\mathcal{L}}$) consists of centerline vectors $\mathcal{L}=\{l_{1},\ldots,l_{m_{l}}\}$ that represent the centerlines. Each centerline $l_{i}$ is transformed into vectors $\mathcal{L}_{i}=\overrightarrow{p_{i}^{1}p_{i}^{2}},\overrightarrow{p_{i}^{2}p_{i}^{3}},\ldots$ based on consecutive points $p_{i}^{j}$. The boundary ($\mathcal{B}$) includes the boundary vectors $\mathcal{B}=\{b_{1},\ldots,b_{m_{b}}\}$, converted similarly to the roads: $b_{j}\to\mathcal{B}_{j}=\overrightarrow{h_{j1},h_{j2}},\overrightarrow{h_{j2}h_{j3}},\ldots$. These vectors are processed by the vector embedding module, which maps each 5D vector $\vec{v}_{i}$ to a high-dimensional feature $F_{\vec{v}_{i}}\in\mathbb{R}^{C}$ via a two-layer MLP. The outputs are aggregated into feature matrices: $F_{road}\in\mathbb{R}^{N_{r}\times C}$, $F_{centerline}\in\mathbb{R}^{N_{l}\times C}$, and $F_{boundary}\in\mathbb{R}^{N_{b}\times C}$, where $N_{(\cdot)}$ denotes the total number of vectors (e.g. $N_{r}=\sum_{j=1}^{m_{r}}\text{length}(\mathcal{R}_{j})$).
@@
Subsequently, the vector tokens are input into a transformer network. MAT consists of stacked MAT blocks to extract hierarchical features, each block containing Path-Aware Attention (PA), Spatial Attention (SA), and feed-forward network (FFN). In the association head, each road token requires the pooling to obtain a representative token $\bar{F}_{road}$ for the current road. The association of the centerline with the road is calculated by combining the attention between $\bar{F}_{road}$ and $F_{centerline}$, generating the probability distribution of the associations of the SD-OP map. The association probabilities are further refined by a post-processing method to enforce topological constraints. The details of the implementation and the module are provided in Appendix F.
@@ 5.2 Path-aware Attention
Path-Aware Attention (PA) is designed to extract locally stable geometric features that preserve the structural arrangement of elements.
Topological Ordering for Efficient Attention. To achieve real-time inference, our framework employs Group Attention, which reduces computational complexity from $O(N^{2})$ to linear $O(N)$ by restricting interactions to local windows. However, this efficiency comes with a constraint: the effectiveness of group attention depends heavily on the semantic meaningfulness of token order. PA addresses this by introducing Topological Ordering as an inductive bias. Unlike random ordering, which would fragment the graph into unrelated segments, PA explicitly models long-range dependencies by constructing paths from root to leaf nodes. This ensures that topologically connected predecessor and successor nodes are placed adjacently in the sequence.
Mechanism. Specifically, we identify all valid paths from a starting point to an endpoint. We then reorder vector tokens to align with these path indices. This strict ordering allows the subsequent grouped attention (with size $k$) to focus solely on topologically relevant neighbors. Finally, tokens are reversed to their original sequence, and features of tokens appearing in multiple paths are averaged.
@@ 5.3 Spatial Attention
While PA captures topological connectivity, it may miss interactions between geometrically adjacent but topologically distant segments (e.g., parallel lanes or disconnected road boundaries). The Spatial Attention (SA) mechanism complements PA by capturing instance-level interactions across a wider spatial scope via Spatial Ordering.
Motivation: Vector Serialization as Geometric Clustering. Similar to PA, SA relies on a specific ordering strategy to maximize the efficacy of group attention. However, instead of following graph connectivity, SA utilizes Vector Serialization to cluster tokens based on geometric proximity. This serves as a spatial inductive bias: it forces physically adjacent entities—even if they belong to different map layers or are disconnected—to be placed adjacently in the 1D sequence. Consequently, when the sequence is sliced into groups, highly correlated tokens naturally fall into the same attention bucket, enabling the model to handle GPS offsets and map alignment robustly.
Mechanism: Attention with Vector Serialization. As illustrated in Fig. 5 green part, we implement this spatial ordering through the following steps:
1) Coordinate Discretization: Each vector token $\vec{v}_{i}$ is encoded into a 3D discrete coordinate $(x,y,r)$, representing the quantized grid location and orientation of the vector.
2) Serialization via Space-Filling Curves: We employ a space-filling curve function $\varphi^{-1}$ (e.g., the Hilbert curve hilbert1935stetige) to map the 3D coordinates to a single 1D index. This mapping is crucial as it preserves spatial locality in the 1D domain better than simple row-major scanning.
3) Grouped Attention & Restoration: Tokens are reordered based on their 1D indices. Self-attention is performed within each group to aggregate spatial context, followed by an inverse operation to restore the original order.
@@ 5.4 Association and Loss Function
Association. The association between the roads and the centerline is calculated through a mechanism of cross-attention. For each road $j$, we first aggregate its token features $\{F^{road}_{j1},\ldots,F^{road}_{jN}\}$ into a representative feature $\bar{F}^{road}_{j}=\frac{1}{N}\sum_{n=1}^{N}F^{road}_{jn}$, where NN denotes the number of road tokens on roads $r_{j}$ and $F^{road}_{jn}\in\mathbb{R}^{d}$. The association probability $Prob_{ij}$ between the centerline $i$ and the road $j$ is then calculated as:
$Prob_{ij}=\exp\left(\frac{F^{cl}_{i}\cdot\bar{F}^{road}_{j}}{\sqrt{d}}\right)/\sum_{k=1}^{K}\exp\left(\frac{F^{cl}_{i}\cdot\bar{F}^{road}_{k}}{\sqrt{d}}\right),$ (4)
where $F^{cl}_{i}\in\mathbb{R}^{d}$ is the token feature of the centerline, $d$ is the dimension of the feature, and $K$ represents the total number of roads. This formulation normalizes the similarity scores in all roads for each centerline $i$, ensuring a valid probability distribution.
Loss Function. Followed by liu2023graphmm; ren2021mtrajrec, we optimize the model using a combination of cross-entropy loss (CE) and connection temporal classification (CTC) loss. The total loss is a weighted sum:
@@ 5.5 Topology Post-Process
We formalize topological decoding as a structured prediction on the entire path of the centerline $\mathcal{P}_{j},\ j\in[1,\cdots,K]$. $K$ is the number of total paths. The two-stage decoding process operates as follows:
@@
Topological-constraint Beam Search. Based on beam search, topological constraint beam search makes the following two improvements: 1. Modify the one-way search to implement a bidirectional search starting at $T_{max}$. 2. When generating new predictions, instead of using the approach of taking the maximum value from all roads, we decode under the constraint of connectivity provided in the road network $\mathcal{E}_{r}$, thus ensuring that the connectivity of the road sequence corresponding to the lane path in the decoding result is consistent with the representation of the road network. Detailed expressions of the topological constraints beam search, including formula descriptions, are included in the appendix F.
@@ 6 Experiment
Table 1: Result on OMA Val set. La. means latency. $M$, PM, GM, MA means map matching, graph matching, point matching and map association method. P-M and M-M means the model is trained and inferred by path-to-map or map-to-map.
Methods Present Type Paradigm Val set
NR-F150:95 La./ms
HMM DBLP:conf/gis/NewsonK09 SIGSPATIAL’09 $M$ P-M 70.1 465
DeepMM feng2020deepmm TMC’20 $M$ P-M 70.6 328
MTrajRec ren2021mtrajrec KDD’21 $M$ P-M 71.2 849
GraphMM liu2023graphmm TKDE’23 $M$ P-M 72.3 469
EAM3 tang2025elevation TITS’25 $M$ P-M 72.9 345
KNN stuart2010matching GM M-M 68.6 299
GMT he2024learnable TPAMI’24 GM M-M 68.8 93
FastMAC zhang2024fastmac CVPR‘24 PM M-M 70.3 62
MAT-T (Ours) MA M-M 78.2 34
MAT-L (Ours) MA M-M 78.7 70
Table 2: Result on OMA Test set. La. means latency. $M$, PM, GM, MA means map matching, graph matching, point matching and map association method. P-M and M-M means the model is trained and inferred by path-to-map or map-to-map. SGG. means SeqGrowGrpah xie2025seqgrowgraph.
Methods Present Type Paradigm NR-F150:95 on Test Sets La./ms
MapTR MapTRv2 SGG.
HMM DBLP:conf/gis/NewsonK09 SIGSPATIAL’09 $M$ P-M 33.1 36.0 46.5 561
DeepMM feng2020deepmm TMC’20 $M$ P-M 32.8 33.5 43.9 733
MTrajRec ren2021mtrajrec KDD’21 $M$ P-M 34.1 36.8 48.9 1593
GraphMM liu2023graphmm TKDE’23 $M$ P-M 35.9 38.7 49.2 889
EAM3 tang2025elevation TITS’25 $M$ P-M 36.3 39.1 50.8 679
KNN stuart2010matching GM M-M 31.7 34.6 43.5 313
GMT he2024learnable TPAMI’24 GM M-M 30.1 33.4 42.3 105
FastMAC zhang2024fastmac CVPR‘24 PM M-M 32.0 35.8 45.0 81
MAT-T (Ours) MA M-M 41.5 44.8 54.8 35
MAT-L (Ours) MA M-M 41.9 45.0 54.9 74
6.1 Implementation Details
We train our models from scratch for a total of 50 epochs using the AdamW optimizer. A cosine-decay learning rate scheduler is employed, incorporating a linear warm-up phase of two epochs. The initial learning rate, weight decay, and batch size are set to $0.0001$, $0.05$, and $128$, respectively. All experiments are conducted on NVIDIA A6000 GPUs. The latency is measured on an NVIDIA A6000 GPU paired with an Intel(R) Xeon(R) Platinum 8369B CPU. Note that MAT-T and MAT-L share identical architectural components and training configurations, differing only in the number of Transformer blocks to balance real-time efficiency and model capacity, which details are shown in Tab. 22 in Appendix.
@@ 6.2 Result
To evaluate the performance of different methods on the OMA dataset, we categorize existing approaches into map matching methods DBLP:conf/gis/NewsonK09; feng2020deepmm; ren2021mtrajrec; liu2023graphmm; tang2025elevation, graph matching methods he2024learnable, and point matching methods zhang2024fastmac. For map matching, we traverse every path on the OP map as a GPS trajectory to match against the SD map. For graph and point matching, we treat the SD and OP maps as separate graphs or point clouds.
Val set. As shown in Table 1, MAT achieves the optimal balance between precision and efficiency on the validation set. Specifically, compared to traditional, Seq2Seq, and graph-based map matching methods, MAT-T demonstrates significant improvements in $\text{NR-F1}^{50:95}$, outperforming HMM DBLP:conf/gis/NewsonK09 by 8.1%, DeepMM feng2020deepmm by 7.6%, MTrajRec ren2021mtrajrec by 7.0%, and the recent EAM3 tang2025elevation by 5.3%, all while maintaining a low inference latency of 34 ms. Moreover, MAT-T achieves superior performance compared to graph matching techniques like GMT he2024learnable and point matching approaches such as FastMAC zhang2024fastmac, proving its effectiveness in handling the heterogeneity between SD and OP maps which typically challenges traditional matching or point cloud algorithms.
Test set. To further verify the model’s robustness against varying noise patterns crucial for real-world deployment, we extended the evaluation on the Test set to include OP maps generated by three distinct methods: MapTR liaomaptr, MapTRv2 maptrv2, and SeqGrowGraph xie2025seqgrowgraph, as summarized in Table 2. The results demonstrate that our proposed MAT method consistently outperforms all state-of-the-art baselines across these disparate generators without requiring specific fine-tuning. On the SeqGrowGraph part, MAT-L achieves an $\text{NR-F1}^{50:95}$ of 54.9%, surpassing the strongest baseline EAM3 tang2025elevation by 4.1%, while on the more challenging MapTR dataset, which exhibits different topological error patterns, MAT-L maintains a significant lead of 5.6% over EAM3. This consistent superiority confirms that MAT’s architecture effectively generalizes across distinct noise distributions—ranging from fragmentation in MapTR to connectivity issues in growing-based methods—thereby validating the strong cross-generator generalization ability of the proposed Online Navigation Refinement framework.
6.3 Ablation Study
The ablation study experiment is conducted with MAT-L in the val set of OMA, with latency measured using a NVIDIA A6000.
Table 3: Ablation study of structure. Post. means post process. Bd. means Boundary. La. means latency.
SA PA Bd. Post. NR-F150:95 La./ms
Baseline (PTv3) 61.8 59
62.1 77
74.1 61
77.8 64
78.5 69
78.7 70
Table 4: Ablation study of loss and pool method. La. means latency.
CE CTC Avg. Max NR-F150:95 La./ms
78.4 70
67.7 70
78.5 70
78.7 70
Table 5: Cross-validation between Boston and Singapore in the OMA dataset
     City      Val      Test
     Singapore $\rightarrow$ Singapore      77.5      49.5
     Boston $\rightarrow$ Singapore      77.3      48.7
     Boston $\rightarrow$ Boston      79.4      42.9
     Singapore $\rightarrow$ Boston      79.0      42.6
Structure. As shown in Tab. 4, ablating path-aware (PA) and spatial attention (SA) reveals that PA+SA achieves the highest $\text{NR-F1}^{50:95}$ (+3.7% vs. PA-only, +15.7% vs. SA-only). PA-only outperforms SA-only (74.1% vs. 62.1%), confirming the critical role of topological awareness. Boundary improve +0.7% with a 5ms latency cost. Post-processing further enhances accuracy (+0.2%) without efficiency trade-offs.
Loss Function and Road Pooling Method. Tab. 4 shows an ablation study on loss functions and road pooling methods. The model exhibits strong robustness, with consistent performance across different settings, indicating low sensitivity to these components. Combining CE and CTC losses improves NR-F1 by +0.3% over CE alone and by +11.0% over CTC alone, while all variants perform competitively. Average pooling outperforms maximum pooling by only +0.2%, further confirming the model’s insensitivity to pooling strategy. These results suggest that the effectiveness of the method is due to its inherently robust design, rather than specific loss or grouping choices.